OpenStudy (anonymous):

please let f(x)= 1/sqroot x-3 . evaluate the function at ( f(x) - f(4) ) / x-4 . simplify your answer.

OpenStudy (anonymous):

\[f(x)=\frac{1}{\sqrt{x-3}}\] \[f(4)=\frac{1}{\sqrt{4-3}}=1\] \[\frac{f(x)-f(4)}{x-4}=\frac{\sqrt{x-3}-1}{x-4}\]

OpenStudy (anonymous):

now as for "simplifying" i believe what they really mean is rationalize the numerator and cancel

OpenStudy (anonymous):

multiply top and bottom by \[\sqrt{x-3}+1\] so get \[\frac{\sqrt{x-3}-1}{x-4}\times \frac{\sqrt{x-3}+1}{\sqrt{x-3}+1}\] \[\frac{x-3-1}{(x-4)(\sqrt{x-3}+1)}\] \[\frac{x-4}{(x-4)(\sqrt{x-3}+1)}=\frac{1}{\sqrt{x-3}+1}\]